The points (p,-8)and (-10,2) fall on a line with a slope of -5.
What is the value of p?


p=


Answer by rothauserc(4718)

slope = (y2 - y1) / (x2 - x1)
:
-5 = (2 - (-8)) / (-10 - p)
:
-5 = 10 / (-10 - p)
:
-5(-10 - p) = 10
:
50 + 5p) = 10
:
5p = -40
:
************************************
p = -8
:
check the answer
:
-5 = (2 - (-8)) / (-10 - (-8)
:
-5 = 10 / -2
:
-5 = -5
:
our answer checks
************************************
: